A car covers 225 km in 2.5 hours. How far can the car do in 1.75 hours?
Answer by harpazo(655)

Let d = distance car can travel
Set up a proportion and solve for d by cross-multiplication method.

225/d = 2.5/1.75

2.5d = 225(1.75)

2.5d = 393.75

d = 393.75/2.5

d = 157.5 km